# MS-Occ: Multi-Stage LiDAR-Camera Fusion for 3D Semantic Occupancy Prediction

### v1

@@ IV-C Main Results
Comprehensive comparisons are conducted between MS-Occ and state-of-the-art approaches across three modalities: camera-only, LiDAR-only, and LiDAR-camera fusion. As quantitatively shown in Table I, the proposed MS-Occ framework achieves superior overall performance with an IoU of 32.1 and an mIoU of 25.3, demonstrating particularly significant gains in vulnerable road user (VRU) categories (pedestrians, bicycles, motorcycles) and small object detection (traffic cones, man-made structures). Compared to the baseline M-CONet [10], these critical classes exhibit substantial IoU improvements of 11.0%, 10.5%, 11.6%, 7.8%, and 8.0%, respectively. These improvements are especially important for autonomous driving safety due to the high kinematic uncertainty of VRUs, the low reflectivity of traffic cones, and the structural regularity requirements for man-made infrastructure.
In comparison with state-of-the-art LiDAR-camera fusion methods, MS-Occ consistently outperforms across different fusion stages. Specifically, it achieves a 2.4% mIoU and a 0.7% IoU improvement over OccLoff [9], a representative middle-stage fusion method, highlighting its competitive cross-modal integration capability. Moreover, compared to Co-Occ, a leading late-stage fusion approach, MS-Occ yields even greater improvements of 3.4% in mIoU and 1.5% in IoU, underscoring its superior performance over late-stage fusion strategies.
Fig. 5 illustrates that MS-Occ produces 3D occupancy predictions with fewer false positives in occluded regions, improved boundary delineation for small-scale objects, and more consistent geometric alignment with sensor observations.
Figure 5: Visualization results on the nuScenes-OpenOccupancy dataset [34, 10]. The leftmost column shows the surround-view images, while the next three columns present 3D semantic occupancy predictions from M-CONet, MS-Occ (ours), and the ground truth annotations, respectively. Please zoom in for finer details.
@@ IV-D Ablation Study
Effect of Fusion Stages. Ablation studies are conducted to evaluate the impact of different fusion stages on the overall model performance, as presented in Table LABEL:tab:ablation_MSOcc_FusionStage. The middle-stage feature fusion effectively compensates for the semantic deficiencies of the LiDAR modality and the geometric limitations of the camera modality. Removing this stage results in a significant performance drop, 4.2% decrease in mIoU. The voxel fusion in the late stage primarily refines feature integration in 3D space. Eliminating this stage leads to a 1.4% decrease in mIoU.
@@
Effect of Main Components. Table LABEL:tab:ablation_MSOCC_Components presents the impact of key components on model performance, validating the effectiveness of the proposed modules, including the Gaussian-Geo module, Semantic-Aware module, AF module, and HCCVF module. The ablation results reveal the following: (1) Removing the Gaussian-Geo module results in moderate performance degradation, with a 0.9% decrease in mIoU; (2) Eliminating the Semantic-Aware module causes a substantial performance drop, reducing mIoU by 3.2%; (3) Discarding the AF module leads to a noticeable performance decline, lowering mIoU by 0.4%; and (4) Removing the HCCVF module results in considerable performance deterioration, decreasing mIoU by 1.0%. These results demonstrate that each component contributes meaningfully to the entire MS-Occ framework.
@@
Ablation of HCCVF. Table IV presents the ablation study on fusion strategies in the late fusion stage, focusing on the proposed HCCVF module. To evaluate its effectiveness, HCCVF is compared with two alternative fusion methods adapted from [15]: (1) Summation-based fusion, where voxel features from both modalities are directly added and passed through an MLP for feature fusion; and (2) Concatenation-based fusion, where voxel features are concatenated along the sequence dimension and subsequently processed by an MLP. The experimental results demonstrate that the HCCVF module achieves superior performance compared to these conventional fusion strategies. Specifically, HCCVF outperforms the summation-based method by 0.8% in mIoU and surpasses the concatenation-based method by 0.6% in mIoU.
TABLE II: Ablation study of fusion stages.
Middle Stage Late Stage IoU mIoU
30.2 21.1
31.4 23.9
32.1 25.3
TABLE III: Ablation study of main components.
Gaussian-Geo Semantic-aware HCCVF AF IoU mIoU
30.6 22.1
31.7 24.4
31.6 24.3
31.9 24.9
32.1 25.3
@@
Ablation of Gaussian Kernel. Table V presents an ablation study on the Gaussian kernel length, $\sigma$, in the Gaussian-Geo module to evaluate its impact on model performance. A significant improvement is observed as $\sigma$ increases from 3 to 6, while further increasing $\sigma$ to 10 yields only marginal gains. This suggests that the influence of kernel size reaches a saturation point during the densification of individual point clouds. Given that larger kernel sizes incur higher memory consumption, $\sigma=6$ is selected to achieve a balance between performance and computational efficiency.
TABLE IV: Ablation study of HCCVF.
Fusion Method SC IoU mIoU
Summation 31.6 24.5
Concatenation 31.7 24.7
Self-Attention(ours) 32.1 25.3
TABLE V: Ablation study of Gaussian Kernel.
Kernel Size SC IoU mIoU
3 31.8 24.7
6 32.1 25.3
10 32.2 25.5
@@ V Conclusions
In this letter, MS-Occ is proposed as a novel multi-stage LiDAR-camera fusion framework for 3D semantic occupancy perception, addressing the key limitations of existing unimodal and fusion-based methods. By integrating middle-stage feature fusion with late-stage voxel fusion, MS-Occ achieves robust geometric and semantic alignment across sensing modalities. The Gaussian-Geo module introduces dense geometric priors from LiDAR to enhance image features, while the Semantic-Aware module injects rich semantic context into LiDAR voxel features via deformable attention. In the late fusion stage, the AF and HCCVF modules further refine voxel representations, resolving semantic ambiguities and improving the quality of occupancy predictions. Extensive experiments on the nuScenes-OpenOccupancy dataset validate the effectiveness of MS-Occ, demonstrating superior performance in fine-grained occupancy reasoning, particularly in safety-critical scenarios involving vulnerable road users and small objects.

### v2

@@ IV-C Main Results
IV-C1 Results on NuScenes Dataset
As quantitatively shown in Table I, the proposed MS-Occ framework achieves superior overall performance with an IoU of 32.1 and an mIoU of 25.3, demonstrating particularly significant gains in vulnerable road users (VRUs) categories (pedestrians, bicycles, motorcycles) and small object detection (traffic cones, man-made structures). Compared to the baseline M-CONet [wang2023openoccupancy], these critical classes exhibit substantial IoU improvements of 11.0%, 10.5%, 11.6%, 7.8%, and 8.0%, respectively. These improvements are especially important for autonomous driving safety due to the high kinematic uncertainty of VRUs and the low reflectivity of traffic cones. By fusing rich visual context, MS-Occ effectively mitigates the point sparsity limitations of LiDAR-only methods, elevating performance on challenging classes, such as pedestrian detection from 9.6% (L-CONet) to a superior 29.0% mIoU.
In comparison with state-of-the-art LiDAR-camera fusion methods, MS-Occ consistently outperforms across different fusion stages. Specifically, it achieves a 2.4% mIoU improvement over OccLoff [occloff], a representative middle-stage fusion method. Moreover, compared to Co-Occ, a leading late-stage fusion approach, MS-Occ yields even greater improvements of 3.4% in mIoU and 1.5% in IoU. Fig. 5 illustrates that MS-Occ produces 3D occupancy predictions with fewer false positives and improved boundary delineation. The color scheme used in Fig. 5 corresponds to the class columns in Table I.
IV-C2 Results on SemanticKITTI
To further validate the generalization capability, we evaluate our model on the SemanticKITTI dataset [semantickitti]. As shown in Table II, MS-Occ establishes a new state-of-the-art mIoU of 24.08, substantially outperforming prior methods such as M-CONet (17.25) and OccLoff (22.62). This result demonstrates that the proposed multi-stage fusion architecture is broadly effective and does not overfit a single dataset. Notably, Table II highlights the strong performance of MS-Occ on challenging small objects (e.g., bicyclist, person, traffic-sign) and geometrically complex structures (e.g., vegetation, trunk, fence).
IV-C3 Efficiency Analysis
Table III summarizes the efficiency analytsis. MS-Occ attains state-of-the-art IoU and mIoU while maintaining competitive inference speed and requiring 18.3% fewer parameters. This reflects a highly favorable trade-off, achieving critical accuracy gains for driving safety within a more compact and parameter-efficient architecture.
@@ IV-D Ablation Study
To validate the effectiveness of the proposed components, we perform a series of ablation studies, all conducted on the nuScenes-OpenOccupancy validation set.
TABLE IV: Ablation studies on fusion stages, key components, and design choices. ”w/o” denotes the removal of a module/stage.
Exp. Mid. Stage Late Stage G-Geo Sem-Aware HCCVF AF Kernel IoU mIoU
A w/o Late - - 6 31.4 23.9
B w/o Middle - - Self-Attention - 30.2 21.1
C w/o GG Self-Attention - 31.7 24.4
D w/o SA Self-Attention 6 30.6 22.1
E w/o HCCVF 6 31.6 24.3
F Self-Attention w/o AF 6 31.9 24.9
G Summation 6 31.6 24.5
H Concatenation 6 31.7 24.7
I Self-Attention 3 31.8 24.7
J Self-Attention 10 32.2 25.5
K Self-Attention 6 32.1 25.3
@@
Effect of Fusion Stages. Ablation studies are conducted to evaluate the impact of different fusion stages on the overall model performance, as presented in Table IV (Exp. A-B). The middle-stage feature fusion effectively compensates for the semantic deficiencies of the LiDAR modality and the geometric limitations of the camera modality. Removing this stage results in a significant performance drop of 4.2% in mIoU. The voxel fusion in the late stage primarily refines feature integration in 3D space. Eliminating this stage leads to a 1.4% decrease in mIoU.
@@
Effect of Main Components. Table IV (Exp. C-F) presents the impact of key components on model performance. The ablation results reveal the following: (1) Removing the Gaussian-Geo module results in a 0.9% decrease in mIoU; (2) Eliminating the Semantic-Aware module causes a substantial 3.2% drop in mIoU; (3) Discarding the AF module leads to a 0.4% decline in mIoU; and (4) Removing the HCCVF module results in a 1.0% decrease in mIoU. These demonstrate that each component contributes meaningfully to MS-Occ framework.
@@
Ablation of HCCVF. Table IV (Exp. G-H) presents the ablation study on fusion strategies in the late fusion stage. To evaluate its effectiveness, HCCVF is compared with two alternative fusion methods adapted from [MultiModalObjectDetection]: (1) Summation-based fusion and (2) Concatenation-based fusion. The experimental results demonstrate that the HCCVF module achieves superior performance compared to these conventional fusion strategies. Specifically, HCCVF outperforms the summation-based method by 0.8% in mIoU and surpasses the concatenation-based method by 0.6% in mIoU.
@@
Ablation of Gaussian Kernel. Table IV (Exp. I-J) presents an ablation study on the Gaussian kernel length, $\sigma$, in the Gaussian-Geo module. A significant improvement is observed as $\sigma$ increases from 3 to 6, while further increasing $\sigma$ to 10 yields only marginal gains. This suggests that the influence of kernel size reaches a saturation point. Given that larger kernel sizes incur higher memory consumption, $\sigma=6$ is selected to achieve the best balance between performance and computational efficiency.
@@ V Conclusions
In this letter, we introduce MS-Occ, a multi-stage LiDAR-camera fusion framework for 3D semantic occupancy. MS-Occ couples middle-stage feature fusion with late-stage voxel fusion to achieve strong geometric and semantic alignment. Gaussian-Geo injects geometric priors into image features, Semantic-Aware enriches LiDAR voxels, and AF and HCCVF refine voxel representations to resolve semantic ambiguities. Experiments on nuScenes and SemanticKITTI confirm the effectiveness of MS-Occ, with notable gains in detecting VRUs and other small, safety-critical objects. Nevertheless, a limitation is that performance gains are less pronounced for large objects. This is likely because strong visual cues already make camera-only methods highly competitive in these cases, whereas our fusion mechanism is mainly designed to enhance fine-grained details and resolve geometric ambiguities in smaller or complex objects. Future work could explore multi-scale fusion to better capture structural information from large objects. Furthermore, the current framework assumes accurate LiDAR-camera extrinsic calibration. While our model shows some robustness, improving its resilience to real-world sensor misalignments is an important future direction.